# IterPref: Focal Preference Learning for Code Generation via Iterative Debugging

### v2

@@ Abstract
Preference learning enhances Code LLMs beyond supervised fine-tuning by leveraging relative quality comparisons. Existing methods construct preference pairs from $n$ candidates based on test case success, treating the higher pass rate sample as positive and the lower as negative. However, this approach does not pinpoint specific errors in the code, which prevents the model from learning more informative error correction patterns, since aligning failing code as a whole lacks the granularity needed to capture meaningful error-resolution relationships. To address these issues, we propose IterPref, a new preference alignment framework that mimics human iterative debugging to refine Code LLMs. IterPref explicitly locates error regions and aligns the corresponding tokens via a tailored DPO algorithm. To generate informative pairs, we introduce the CodeFlow dataset, where samples are iteratively refined until passing tests, with modifications capturing error corrections. Extensive experiments show that a diverse suite of Code LLMs equipped with IterPref achieves significant performance gains in code generation and improves on challenging tasks like BigCodeBench. In-depth analysis reveals that IterPref yields fewer errors. Our code and data will be made available recently.
IterPref: Focal Preference Learning for Code Generation
via Iterative Debugging
@@ 1 Introduction
Figure 1: IterPref achieves significant performance gains over DPO variants on challenging coding tasks, BigCodeBench-Hard, with advanced code LLMs Deepseek-Coder-7B-Instruct and Qwen2.5-Coder-7B.
Preference learning offers a promising complement to supervised fine-tuning (SFT) Zhang et al. (2023) for improving code generation accuracy in coding large language models (Code LLMs). A major challenge lies in the scarcity of high-quality preference data. Existing methods Zhang et al. (2024, 2025); Liu et al. (2024c) mainly rely on unit test feedback to construct preference pairs. In these approaches, a Code LLM generates multiple code snippets as candidates and evaluates each against a suite of test cases. The snippet with the higher pass rate is considered preferred, while the one with the lower pass rate is marked as dispreferred, which forms the pair for preference learning such as Direct Preference Optimization (DPO) Rafailov et al. (2023). However, synthesizing preference pairs based solely on pass rate has notable limitations. Specifically, a snippet with a low pass rate may only require minor modifications to become correct, as errors may be isolated to specific parts of the code, as shown in Figure 2. Relying on full preference learning in this context can introduce noise during optimization Pal et al. (2024); Chen et al. (2024); Wu et al. (2024). It not only hinders the model from learning more effective error correction patterns but also increases the risk of overfitting. The underlying issue is that such preference pairs cannot explicitly identify which parts of the code need to be aligned.
@@
To tackle this challenge, we draw inspiration from the way developers debug code. Typically, a programmer first locates the module that generates errors based on execution feedback and then focuses on fixing that specific portion. Following this human approach, we introduce IterPref, a novel framework for preference learning in Code LLMs that leverages iterative debugging insights. Rather than only using pass rate to measure the degree of preference, IterPref derives preference pairs from debugging process itself. By contrasting critical tokens between a corrected version and its preceding iteration explicitly, the framework guides the model to learn fine-grained alignment for key errors.
Specifically, we first propose a new iterative debugging dataset, termed CodeFlow, to generate preference pairs. In CodeFlow, a Coding LLM initially generates a function-level code snippet along with its test case, then refines the snippet iteratively until it passes all unit tests. Preference pairs are formed by treating the final correct version as preferred and earlier failed versions as dispreferred. These pairs allow for the efficient annotation of critical tokens that distinguish the correct code from its erroneous predecessors, thus highlighting the key changes required for pairwise optimization.
For effective preference learning, we design an improved DPO algorithm to explicitly mark the tokens to be optimized from dispreferred samples. In our approach, the preferred sample receives rewards for all tokens to promote a full understanding of the function’s structure, while only the error-specific tokens in dispreferred samples (labeled in CodeFlow) are penalized. This targeted strategy works collectively with the CodeFlow dataset, minimizing noise from correct code and sharpening the mode’s focus on learning error patterns.
We conduct extensive experiments to validate the effectiveness of IterPref. With only 59k preference pairs, IterPref achieves significant performance gains on HumanEval and MBPP across various base and instruct-tuned Code LLMs. Additionally, as shown in Figure 1, IterPref demonstrates superior results on complex coding tasks, like BigCodeBench. Our contributions are as follows:
@@
    We leverage the idea of iterative debugging to tackle the challenges in Code LLM preference learning, enabling more focused alignment on critical error tokens.
@@
    We construct a new function-level dataset that iteratively tracks token differences across preference pairs, and propose a tailored adaptation of the DPO algorithm that avoids unnecessary optimization noise.
@@
    IterPref consistently improves performance across diverse benchmarks and various base and instruct-tuned Code LLMs.
2 Related Work
Code Language Models. Powerful Code LLMs like Qwen2.5-Coder Hui et al. (2024a), DeepSeek-Coder Guo et al. (2024), StarCoder Li et al. (2023); Lozhkov et al. (2024), Magicoder Wei et al. (2024) and EpiCoder Wang et al. (2025b) demonstrate their capabilities in various code generation tasks, such as Humaneval Chen et al. (2021), MBPP Austin et al. (2021), BigCodebench Zhuo et al. (2025), and LiveCodebench Jain et al. (2024). Current Code LLMs primarily focus on supervised fine-tuning during the post-training stage. While SFT enables Code LLMs to learn the correct patterns, it fails to effectively make them aware of incorrect patterns or how to rectify errors in code. In this work, IterPref framework aims to enable Code LLMs to further learn through pairwise contrasting of critical tokens Lin et al. (2024), allowing Code LLMs to continually improve.
Code Preference Construction. PLUM Zhang et al. (2024) introduced preference pair construction by ranking candidate code solutions based on test case passed. Code-Optimise Gee et al. (2025) further integrates efficiency as another learning signal, augmented with annotations from unit test feedback and execution time. CodeDPO Zhang et al. (2025) also targets correctness and efficiency, proposing an improved PageRank-inspired iterative algorithm for more precise pair selection. AceCoder Li et al. (2024) selects pairs with distinct pass rate difference. Meanwhile, DSTC Liu et al. (2024b) constructs preference pairs using self-generated code snippets and tests, without relying on external LLMs for generation and annotation.
3 IterPref Framework
The IterPref framework mimics human iterative debugging to refine Code LLMs. It explicitly identifies error regions and focuses on aligning the corresponding tokens through a tailored DPO algorithm. To achieve this, IterPref follows two steps: 1) synthesizing preference code pairs through an iterative debugging process and locate error regions within code (Section 3.1), and 2) performing fine-grained and focal alignment by contrasting critical tokens via the designed DPO algorithm (Section 3.2).
@@ 3.1 Synthesize Preference Code Snippets
In contrast to previous methods that synthesize preference pairs based only on pass rate, IterPref synthesizes preference code snippets from the iterative debugging process. In this process, an initial code snippet is refined until it passes the test cases, and a preference pair is constructed between the final correct version and the previous iteration.
Generate Raw Code Snippets and Tests
To obtain diverse and complex code data, we adopt the prompt template from EpiCoder Wang et al. (2025a), leveraging its feature tree-based synthesis framework to generate high-quality code and test cases with GPT-4o OpenAI (2024). Using this approach, the LLM generates a coding task instruction, the corresponding code snippet, a test file, and a suitable execution environment.
Iterative Refinement through Verification
Relying solely on generated code and test cases is unreliable Ma et al. (2025), as LLMs cannot guarantee the correctness of generated code. Therefore, we verify each code and refine incorrect versions through an iterative debugging process.
For each code and its corresponding test file, we evaluate the correctness of the code through unit test verification (e.g., using assertions). As shown in Figure 4, when the code fails the unit test or encounters runtime errors, we collect the error information and refine the code iteratively until it passes the test. The pass rate at the $T$-th iteration is reported in Figure 3. At the first attempt (iter0), only 36.7% of the code samples pass their corresponding test file, indicating that debugging is necessary for the remaining code. Failed codes go through continual refinement, with the pass rate gradually approaching 67.5%. The pass rate rises sharply from iter 1 to iter 3 and then slows. Between iter 4 and iter 5, only 1.2% of cases improve, indicating incremental benefits from further iterations. Thus, additional iterations are not considered.
Figure 4: Overview of IterPref. IterPref constructs preference pairs via iterative debugging, treating the correct version as preferred and the previous as dispreferred. DPO adaptations enable code LLMs to learn the correct pattern from the preferred code while highlighting critical tokens with a masking strategy in the dispreferred sample.
Our goal is to collect the program changes made during the iterative debugging process. Therefore, codes that succeed on the first iteration, along with those that cannot be corrected within 5 iterations, are discarded. This iterative debugging process mimics how humans write programs by constantly modifying the code based on test results. The key code changes during the iteration serve as promising reward signals that guide the code from incorrect to correct. Driven by the reward signals from this iterative process, we treat the final correct code as the chosen sample and randomly select an earlier version of the code as the rejected sample, forming the pair ($y^{+}$, $y^{-}$) for preference-based learning.
Critical Difference Extraction
@@
4:  $\mathcal{D}^{+}=\{l^{+}\in y^{+}_{\text{lines}}\mid l^{+}\notin\text{LCS}\}$
5:  $\mathcal{D}^{-}=\{l^{-}\in y^{-}_{\text{lines}}\mid l^{-}\notin\text{LCS}\}$
@@
For each ($y^{+}$, $y^{-}$) pair, we extract the critical differences that drive the functional outcome. Specifically, we identify the sets of difference lines $D^{+}$ and $D^{-}$ based on the Longest Common Subsequence (LCS), where the lines not belonging to the LCS are considered as the difference lines, as shown in Algorithm 1. The extraction process does not require annotations from humans or LLMs. After extraction, the key modifications between pairs are constrained within $D^{+}$ and $D^{-}$. $D^{+}$ and $D^{-}$ record the changes made between the preferred and dispreferred versions, which are responsible for the functional differences between the pairs.
Quality Control for Preference Pairs
We synthesize 104k instruction data through iterative refinement and implement several measures to ensure the quality of the synthesized data. In particular, we apply rule-based filtering to remove trivial or uninformative samples, including: (i) $D^{-}$ consists only of comments; (ii) $D^{-}$ exceeds 20 lines; (iii) $y^{+}$ or $y^{-}$ exceeds 2048 tokens; and (iv) Samples where the abstract syntax tree (AST) of $y^{+}$ and $y^{-}$ are identical.
After applying these filters, 84k samples remained. Next, for each sample, we employ the LLM-as-a-judge methodology using GPT-4o to assess whether there is a significant logical difference between $y^{+}$ and $y^{-}$. We further filter out pairs where the differences are limited to code formatting, comments, variable names, whitespace, or blank lines. These efforts ensure that the selected and rejected samples reflect key functional differences. The final dataset consists of 59,302 samples, ready for preference-based alignment training.
3.2 Fine-grained Alignment
@@
DPO has proven effective in multiple domains like mathematics Lai et al. (2024). However, the objective in Eq. (2) may not be fully suitable for preference-based alignment in code generation, as a large portion of the tokens in $y^{+}$ and $y^{-}$ are identical, with only minor differences. This can confuse the policy model in identifying the critical differences necessary for functional correctness, and diminish alignment gains Pal et al. (2024); Chen et al. (2024); Wu et al. (2024).
@@
$K^{-}=\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)}&=\log\frac{\prod_{i=1}^{L}\pi_{\theta}(y_{i}^{-}|x)}{\prod_{i=1}^{L}\pi_{\text{ref}}(y_{i}^{-}|x)}\\ &=\sum_{i=1}^{L}\log\frac{\pi_{\theta}(y_{i}^{-}|x)}{\pi_{\text{ref}}(y_{i}^{-}|x)}\end{aligned}$
@@
Eq. (4) encourages the code LLMs to learn correct code generation patterns from $y^{+}$. Eq. (5) explicitly focuses on contrasting critical tokens in $y^{-}$ while keeping the correct tokens in the dispreferred code uninvolved in the loss computation through masking the tokens that do not appear in $D^{-}$. By penalizing critical tokens responsible for functional errors and avoiding over-optimization on tokens common to both $y^{+}$ and $y^{-}$, IterPref achieves more fine-grained alignment suitable for code, as opposed to previous sample-level optimization. This helps the code LLMs internalize correct coding patterns and more effectively recognize key token mistakes.
@@
RPO 53.0 45.7 63.0 42.6 28.7 17.2 9.05 6.01 13.1 30.9
Code-Optimise 61.0 54.9 66.5 53.4 31.8 18.8 6.76 6.08 14.9 34.9
IterPref-RPO 73.2 65.2 65.9 53.4 40.3 38.8 18.9 18.2 19.4 43.7
@@ 4 Experiments
Experiment Setup Each instruction-tuned code LLM is fine-tuned for 2 epochs, while base models are fine-tuned for 5 epochs using full-parameter fine-tuning. The model with the lowest validation loss is selected for evaluation. For our IterPref, the learning rate is set to 1e-5 for the 7B code LLMs and 5e-6 for the 14B models, using a global batch size of 128, with a cosine scheduler and warm-up. The maximum sequence length is set to 2048 tokens. For the DPO algorithm, $\beta$ is set to 0.1, and $\alpha$ is set to 1.0 for RPO. $\pi_{\theta}$ and $\pi_{\text{ref}}$ are both initialized with the weights of the evaluated model, while $\pi_{\text{ref}}$ keeps frozen during training.
Benchmarks
We evaluate the Code LLMs using five benchmarks: HumanEval Base Chen et al. (2021), HumanEval Plus Liu et al. (2023), Mostly Basic Python Problems (MBPP Base Austin et al. (2021), MBPP Plus), LiveCodeBench (LCB) Jain et al. (2024) (v5 with problems released between May 2023 and Jan 2025), and Big-CodeBench (BCB) Zhuo et al. (2025). BCB consists of two splits: the instruct split, which involves only instructions, and the complete split, which uses structured docstrings. The hard subset in BCB contains the most challenging fraction of the full set. We use greedy decoding for code generation, with results reported as the pass@1 score. The results of un-tuned Code LLMs are taken from the corresponding leaderboards [^1] https://bigcode-bench.github.io/  [^2] https://evalplus.github.io/ for reference, while other settings are trained by us.
Evaluated Models and Baselines
Evaluated models include DeepSeek-Coder-7B-Instruct-v1.5 Guo et al. (2024), CodeQwen1.5-7B-Chat Bai et al. (2023), and the base models Qwen2.5-Coder-7B Hui et al. (2024b), StarCoder2-15B Lozhkov et al. (2024). The open-source work Code-Optimise Gee et al. (2025) is reproduced using GPT-4o, with 100 solutions sampled at a temperature of 0.6 for each problem. The $DPO_{PvF}$ setting results are reported, with instruct-tuned model appilied DPO, and base models using RPO.

### v3

# Teaching Your Models to Understand Code via Focal Preference Alignment

@@ Abstract
Preference learning extends the performance of Code LLMs beyond traditional supervised fine-tuning by leveraging relative quality comparisons. In existing approaches, a set of n candidate solutions is evaluated based on test case success rates, with the candidate demonstrating a higher pass rate being labeled as positive and its counterpart with a lower pass rate as negative. However, because this approach aligns entire failing code blocks rather than pinpointing specific errors, it lacks the granularity necessary to capture meaningful error-correction relationships. As a result, the model is unable to learn more informative error-correction patterns. To address these issues, we propose Target-DPO, a new preference alignment framework that mimics human iterative debugging to refine Code LLMs. Target-DPO explicitly locates error regions and aligns the corresponding tokens via a tailored DPO algorithm. To facilitate it, we introduce the CodeFlow dataset, where samples are iteratively refined until passing tests, with modifications capturing error corrections. Extensive experiments show that a diverse suite of Code LLMs equipped with Target-DPO achieves significant performance gains in code generation and improves on challenging tasks like BigCodeBench. In-depth analysis reveals that Target-DPO yields fewer errors. Code, model and datasets are in: https://github.com/JieWu02/Target-DPO.
Teaching Your Models to Understand Code via Focal Preference Alignment
Jie Wuthanks: Equal contribution. Work done during the internships of Jie Wu, Haoling Li, and Jianwen Luo at Microsoft Research. Email: {wujie24,li-hl23}@mails.tsinghua.edu.cn.ϕ, Haoling Li⋆ϕ, Xin Zhangthanks: Project leader.⋆π, Xiao Liuπ, Yangyu Huangπ, Jianwen Luoσ, Yizhen Zhangϕ, Zuchao Li, Ruihang Chuthanks: Corresponding author.ϕ, Yujiu Yangϕ, Scarlett Liπ ϕTsinghua University  πMicrosoft Research  σCASIA  Wuhan University
@@ 1 Introduction
Figure 1: Target-DPO achieves significant performance gains over DPO variants on challenging coding tasks, i.e., BigCodeBench-Hard, with Qwen2.5-Coder-7B.
Preference learning offers a promising complement to supervised fine-tuning (SFT) Zhang et al. (2023) for improving code generation accuracy in coding large language models (Code LLMs). Existing methods Zhang et al. (2024, 2025); Liu et al. (2024c) mainly rely on unit test feedback to construct preference pairs. In these approaches, a Code LLM generates multiple code snippets as candidates and evaluates each against a suite of test cases. The snippet with the higher pass rate is considered preferred, while the one with the lower pass rate is marked as dispreferred, which forms the pair for preference learning such as Direct Preference Optimization (DPO) Rafailov et al. (2023). However, this paradigm suffers from two critical drawbacks. First, constructing pairs purely on pass rate cannot guarantee high-quality labels. A high-pass-rate snippet may still carry subtle but crucial bugs, while a low-pass-rate snippet might need only a few modifications to become correct (as shown in Figure 2), resulting in noisy preference data. Second, as errors may be isolated to specific code parts, aligning entire snippets can dilute the correct signal. It forces the model to adjust irrelevant tokens and hinders its ability to learn more specific error patterns Pal et al. (2024); Chen et al. (2024); Wu et al. (2024), which would increase the overfitting risk. The limitations call for a better framework that can pinpoint error regions and apply targeted learning to correct those precise areas.
@@
To tackle these challenges, we draw inspiration from how developers debug code. Typically, a programmer first locates the module that generates errors based on execution feedback and then focuses on fixing that specific portion until all tests pass. Following this human approach, we introduce Target-DPO, a novel framework for preference learning in Code LLMs that leverages iterative debugging insights. Rather than only using pass rate to measure the degree of preference, Target-DPO derives preference pairs from debugging process itself, where the refine steps yield a preference pair, labeling the corrected snippet as preferred and its previous version as dispreferred. By explicitly contrasting the tokens that resolve the error, Target-DPO trains Code LLMs to learn fine-grained alignment for precise error correction, enabling models to truly understand the code.
To support this framework, we efficiently synthesize high-quality preference pairs to create CodeFlow, a novel dataset systematically recording code iterations and corresponding error corrections. Compared to sampling-based methods, CodeFlow enables the efficient creation of preference pairs by (1) generating code snippets and test cases, (2) iteratively refining code until all tests pass, and (3) annotating key token changes between failed and corrected versions. This process ensures that preference learning focuses on the actual error-resolution steps taken by developers.
Building on CodeFlow, we propose an improved DPO algorithm that rewards correct code tokens while penalizing only error-specific tokens in dispreferred samples, minimizing irrelevant noise during preference learning and thus improving efficiency. Comprehensive ablation studies show how to best select dispreferred samples and how much context to include during alignment, verifying our optimal design for code correction.
We conduct extensive experiments on five public datasets to validate the effectiveness of Target-DPO. With only 59k preference pairs, Target-DPO achieves significant performance gains across various base and instruct-tuned Code LLMs. Notably, as shown in Figure 1, Target-DPO attains superior results on complex coding tasks like BigCodeBench. Through detailed ablation studies, we also demonstrate that Target-DPO outperforms alternative strategies by a clear margin. Our contributions are as follows:
@@
    We propose Target-DPO, a novel framework that leverages the idea of iterative debugging to address challenges in preference learning, enabling more precise alignment on critical error tokens.
@@
    We construct a new function-level dataset CodeFlow that iteratively tracks token differences across preference pairs, and propose a tailored adaptation of the DPO algorithm that avoids unnecessary optimization noise.
@@
    Target-DPO consistently improves performance across diverse benchmarks and various base and instruct-tuned Code LLMs.
2 Target-DPO Framework
The Target-DPO framework mimics human iterative debugging to refine Code LLMs. It explicitly identifies error regions and focuses on aligning the corresponding tokens through a tailored DPO algorithm. To achieve this, Target-DPO follows two steps: (1) synthesizing preference code pairs through an iterative debugging process and locate error regions within code, resulting in the creation of CodeFlow (Section 2.1), and (2) performing fine-grained and focal alignment by contrasting critical tokens via the designed DPO algorithm (Section 2.2). The overall workflow is illustrated in Figure 3. In the following sections, we will provide a detailed description of each step.
@@ 2.1 Synthesize Preference Code Snippets
As part of our method, we synthesize 59k preference pairs to enable targeted alignment. In contrast to previous methods that synthesize preference pairs based on pass rate, Target-DPO synthesizes preference code snippets from an iterative debugging process. In this process, an initial code snippet is refined until it passes the test cases, and a preference pair is constructed between the final correct version and the previous iteration.
Generate Raw Code Snippets and Tests. To obtain diverse and complex code data, we adopt the practice of EpiCoder Wang et al. (2025a), utilizing its feature tree-based synthesis framework with GPT-4o OpenAI (2024) to generate high-quality code and test cases. This approach directs the LLM to produce a coding task instruction, the corresponding code snippet, and multiple test cases.
To ensure the quality of the generated test cases, we applied several validation measures, including coverage analysis, LLM-based evaluation, and human verification. Detailed discussions and prompting examples are provided in Appendix A.2.
Iterative Refinement via Verification. LLMs cannot guarantee the correctness of generated code Ma et al. (2025). Therefore, we verify each code sample and refine it through iterative debugging based on execution feedback from verified test cases. As shown in Figure 3, when the initial code fails the unit tests, we collect the error information and refine the code iteratively until it passes the test. The pass rate at the $T$-th iteration is reported in Figure 6 of Appendix A.1.
Figure 3: Method Overview. Target-DPO constructs preference pairs via iterative debugging, treating the correct version as preferred and the previous as dispreferred. DPO adaptations enable code LLMs to learn the correct pattern from the preferred code while highlighting critical tokens with a masking strategy in the dispreferred sample.
Our goal is to collect program changes made during the iterative debugging process. To reduce costs, we discard code that fails to be corrected within five iterations. Although this filter removes some extremely challenging cases, it does not make the generated dataset predominantly easier, as clarified in Appendix A.1. Samples requiring more than five iterations for a solution generally fail to pass all test cases anyway, regardless of additional sampling efforts.
After iterative debugging, we treat the final correct code as the preferred sample and randomly select an earlier version as the dispreferred sample, forming the pair ($^{+}, yy^{-}$) for preference learning.
@@
[…]
@@
Critical Difference Extraction. To identify the targeted regions for alignment, we extract the critical differences responsible for the functional divergence between each ($^{+}, yy^{-}$) pair. Specifically, we pinpoint the sets of differing lines, $D^{+}$ and $D^{-}$, by computing the Longest Common Subsequence (LCS). Lines not part of the LCS are considered difference lines, as detailed in Algorithm 1. Consequently, the key modifications distinguishing the preferred from the dispreferred samples are localized within $D^{+}$ and $D^{-}$. These segments encapsulate the changes driving the functional distinctions and represent the critical regions that Target-DPO aims to contrast and align.
Quality Control for Preference Pairs. We initially synthesize 104k instruction data points through iterative refinement. To further ensure data quality, we implement several filtering measures. Specifically, rule-based filtering is applied to remove trivial or uninformative samples, such as those where: (i) $D^{-}$ consists only of comments; (ii) $D^{-}$ exceeds 20 lines; (iii) $y^{+}$ or $y^{-}$ exceeds 2048 tokens; and (iv) Samples where the abstract syntax tree (AST) of $y^{+}$ and $y^{-}$ are identical.
The application of these filters reduced the dataset to 84k samples. In the next stage, we utilized GPT-4o as an LLM-judge to assess whether a significant logical distinction existed between $y^{+}$ and $y^{-}$. We further filter out pairs where the differences are limited to code formatting, comments, variable names, whitespace, or blank lines. These efforts ensure that the selected and rejected samples reflect key functional differences. The final dataset, consisting of 59k samples, is thus prepared for preference-based alignment training.
2.2 Targeted Preference Alignment
@@
While Direct Preference Optimization (DPO) has demonstrated effectiveness in domains such as mathematics Lai et al. (2024), its standard objective function, as shown in Equation (2), may be suboptimal for preference-based alignment in code generation, as a large portion of the tokens in $y^{+}$ and $y^{-}$ are identical, with only minor differences. This can confuse the policy model in identifying the critical differences necessary for functional correctness, and diminish alignment gains Pal et al. (2024); Chen et al. (2024); Wu et al. (2024).
@@
$\begin{aligned} K^{-}=\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)}&=\log\frac{\prod_{i=1}^{L}\pi_{\theta}(y_{i}^{-}|x)}{\prod_{i=1}^{L}\pi_{\text{ref}}(y_{i}^{-}|x)}\\ &=\sum_{i=1}^{L}\log\frac{\pi_{\theta}(y_{i}^{-}|x)}{\pi_{\text{ref}}(y_{i}^{-}|x)}\end{aligned}$
@@
Equation (4) guides the code LLM to learn correct code generation patterns from $y^{+}$. In contrast, Equation (5) explicitly focuses on contrasting critical tokens within $y^{-}$. It achieves this by masking tokens in the dispreferred code that do not appear in $D^{-}$, thereby excluding correct tokens in $y^{-}$ from the loss computation, as illustrated on the right side of Figure 3. By penalizing critical tokens that cause functional errors and preventing over-optimization on tokens common to both $y^{+}$ and $y^{-}$, Target-DPO achieves a more fine-grained alignment tailored for code, improving upon previous sample-level optimization approaches. This refined strategy enables code LLMs to better internalize correct coding patterns and more effectively identify crucial token-level errors.
@@
DPO 51.8 45.1 63.8 42.9 27.3 16.2 8.1 5.4 12.7 30.4
RPO 53.0 45.7 63.0 42.6 28.7 17.2 9.1 6.0 13.1 30.9
Code-Optimise 61.0 54.9 66.5 53.4 31.8 18.8 6.8 6.1 14.9 34.9
Target-DPO 70.7 64.6 67.2 54.5 39.7 37.7 17.6 16.9 18.7 43.1
Target-RPO 73.2 65.2 65.9 53.4 40.3 38.8 18.9 18.2 19.4 43.7
@@ 3 Experiments
Experiment Setup. For our Target-DPO, the learning rate is set to 1e-5 for the 7B code LLMs and 5e-6 for the 15B models, using a global batch size of 128, with a cosine scheduler and warm-up. The maximum sequence length is set to 2048 tokens. Detailed training settings are presented in Appendix A.3 For the DPO algorithm, $\beta$ is set to 0.1, and for RPO, $\alpha$ is set to 1.0. The rationale behind the choice of $\alpha$ and $\beta$ is supported by ablation studies presented in Appendix C.2. $\pi_{\theta}$ and $\pi_{\text{ref}}$ are both initialized with the weights of the evaluated model, while $\pi_{\text{ref}}$ keeps frozen during training.
Benchmarks.
We evaluate the Code LLMs using multiple benchmarks: HumanEval Base Chen et al. (2021), HumanEval Plus Liu et al. (2023), Mostly Basic Python Problems (MBPP Base Austin et al. (2021), MBPP Plus), LiveCodeBench (LCB) Jain et al. (2024) (v5 with problems released between May 2023 and Jan 2025), and Big-CodeBench (BCB) Zhuo et al. (2025) with instruct and completion splits. We report the pass@1 score under greedy decoding.
Evaluated Models and Baselines.
We evaluate models including DeepSeek-Coder-7B-Instruct-v1.5 Guo et al. (2024), CodeQwen1.5-7B-Chat Bai et al. (2023), as well as base models such as Qwen2.5-Coder-7B Hui et al. (2024b) and StarCoder2-15B Lozhkov et al. (2024). Results for the 32B model are in Appendix C.1. CodeDPO and PLUM are compared using their reported results, as their data and code are currently unavailable. Code-Optimise Gee et al. (2025) is reproduced using GPT-4o, with 100 solutions sampled at a temperature of 0.6 for each problem. The DPO-PvF setting results for Code-Optimise are reported.